• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 15:54:13 on November 26, 2008

    Entry number 250933

    keyword=Day shift summary

    Day Shift Summary

    8:00 Continue reference cell running (3He) from last shift. 8:40 ESR glitch drops lead flows on HRS - HRS crashes. 9:26 ESR still having trouble - go to Controlled Access. Controlled Access activities: 1. HRS recovery (Techs) 2. Beamline HV trouble-shooting (E.C.) 3. Target heater fix. (Jin and Yi) 4. T3 signal into L-HRS scaler (Kalyan) 11:36 Everyone out of hall - cycling HRS and ask for beam permit. 11:50 HRS ready ask for beam. 12:45 Compton ion chamber calibration. 13:05 Back taking data on refernce cell (3He). 14:05 Done with reference cell data - moving to polarized 3He. 14:20 Raster size scan looks good. Ready for production. 16:00 Finish shift in production mode.